# A Survey on Mixture of Experts

### v1

@@ 4.1.1. Sparse
The sparse gating functions activate a selected subset of experts for processing each individual input token, which can be considered as a form of conditional computation (Bengio et al., 2013; Davis and Arel, 2013; Almahairi et al., 2016). The gating functions have been studied extensively, which may be trained by various forms of reinforcement learning and back-propagation, making binary or sparse and continuous, stochastic or deterministic gating decisions (Eigen et al., 2013; Bengio et al., 2015; Rosenbaum et al., 2017, 2019; Clark et al., 2022). Shazeer et al. (Shazeer et al., 2017) pioneered a differentiable heuristic with auxiliary load balancing losses, in which the outputs from expert computations are weighted by their selection probabilities. This introduces a differentiable aspect to the gating process, thereby facilitating the derivation of gradients that can guide the gating function’s optimization. This paradigm has subsequently become predominant in the realm of MoE research. Due to its selection of experts for each input token, this method can be recognized as a gating function with token choice.
@@
Token-Choice Gating. Shazeer et al. (Shazeer et al., 2017) posited the necessity of gating inputs to the top- experts, with , to enhance the efficacy of MoE. The rationale behind this approach is that by simultaneously consulting multiple experts for a given input, the network can effectively weigh and integrate their respective contributions, thereby improving performance. To accommodate the scalability to thousands of experts within a MoE layer, they employ a two-level hierarchical MoE to reduce the branching factor in the context of a large expert count. Subsequent research has largely affirmed that increasing the value of k enhances performance, which has led to the widespread adoption of this top- strategy with . Notwithstanding, the Switch Transformer model (Fedus et al., 2022b) has shown that a top-1 gating strategy (as illustrated in Figure 4 (a)) can also yield competitive results, a finding that has been substantiated and adopted by later studies (Clark et al., 2022). Furthermore, M6-t (Yang et al., 2021) proposed a novel variation of the top-1 gating called expert prototyping, which organizes experts into k groups and then applies top-1 gating in each group. Their experimental results show the training and downstream perplexity of a 16-layer model in order of best to worst: expert prototyping with 4 top-1 gating, 1 top-4 gating, 1 top-16 gating, 1 top-1 gating.
@@
Auxiliary Loss for Token-Choice Gating. Token-choice gating algorithms frequently incorporate an auxiliary loss during training to promote equitable token distribution across experts. Table 1 shows prevalent auxiliary loss functions leveraged in the field. Shazeer et al. (Shazeer et al., 2017) quantify the importance of an expert in relation to a training batch via the batchwise sum of the gate values for that expert. They define an additional loss , added to the overall loss function for the model. This loss, which is equal to the square of the coefficient of variation of the set of importance values and multiplied by a hand-tuned scaling factor , encourages all experts to have equal importance. Although promotes balance in importance, it does not guarantee an even distribution of training examples among experts, which can lead to execution inefficiencies on distributed computing environments. To address this, they introduce a second loss to ensure balanced loads. Building on this foundation, GShard (Lepikhin et al., 2020) define a new differentiable auxiliary loss using a differentiable approximation (the dot-product of mean gates and mean gating decisions per expert), as detailed in Section 2.2. Switch Transformers (Fedus et al., 2022b) and many other subsequent studies (Du et al., 2022; Jiang et al., 2024a; Databricks, 2024; Lieber et al., 2024) have embraced this design, and enhancements (Dai et al., 2024; DeepSeek-AI, 2024; Tianwen Wei, 2024) have been made to cater to diverse requirements. Nevertheless, ST-MoE (Zoph et al., 2022) identified limitations with , particularly at larger scales, leading to unreliable training outcomes. To mitigate this, it introduces the integration of router z-loss , improving training stability without quality degradation by penalizing large logits entering the gating network. Since this loss encourages absolute magnitude of values to be smaller, roundoff errors are reduced, which can be quite impactful for exponential functions such as the gating. Additionally, Mod-Squad (Chen et al., 2023a) posits the difficulty of training multi-task models under such an expert-balancing loss, which may inadvertently force experts to set parameters on conflicting tasks or hinder the potential synergies from parameter sharing across complementary tasks. Instead, it proposes to maximize the mutual information (MI) between experts and tasks to build task-expert alignment. Differently, ModuleFormer (Shen et al., 2023c) proposes to maximize the Mutual Information between experts and tokens. Furthermore, DS-MoE (Pan et al., 2024) extends the application of , calibrating different weightings , in Mixture-of-Attention (MoA, as illustrated in Figure 5 (a)) and FFN MoE modules of different size models.
| Reference | Auxiliary Loss | Coefficient |
| Shazeer et al.(Shazeer et al., 2017), V-MoE(Riquelme et al., 2021) | , | |
| GShard(Lepikhin et al., 2020), Switch-T(Fedus et al., 2022b), GLaM(Du et al., 2022), Mixtral-8x7B(Jiang et al., 2024a), DBRX(Databricks, 2024), Jamba(Lieber et al., 2024), DeepSeekMoE(Dai et al., 2024), DeepSeek-V2(DeepSeek-AI, 2024), Skywork-MoE(Tianwen Wei, 2024) | ||
| ST-MoE(Zoph et al., 2022), OpenMoE(Xue et al., 2024), MoA(Zhang et al., 2022b), JetMoE (Shen et al., 2024) | , | |
| Mod-Squad(Chen et al., 2023a), Moduleformer(Shen et al., 2023c), DS-MoE(Pan et al., 2024) |
@@
Expert Capacity for Token-Choice Gating. In conjunction with load balancing via auxiliary loss, GShard (Lepikhin et al., 2020) incorporates an expert capacity limit, defining a threshold for the number of tokens an expert can process. This can lead to token overflow, where excess tokens are not processed by the designated expert. GShard also proposes a random routing mechanism that selects a secondary expert with a probability proportional to its weight, under the intuition that the contribution of a secondary expert can be negligible, given that the output is a weighted average and the secondary weight is typically small. For the task of image classification with Vision Transformer (ViT) models, Riquelme et al. (Riquelme et al., 2021) enhance the top- gating strategy with Batch Prioritized Routing (BPR), which assigns priority based on higher gating scores rather than the sequence order of tokens. Zoph et al. (Zoph et al., 2022) have demonstrated the efficacy of BPR in the context of MoE language models. Kim et al. (Kim et al., 2021) suggest randomizing token prioritization within sequences to mitigate routing bias towards early-positioned tokens. OpenMoE (Xue et al., 2024) provides a comprehensive analysis of gating mechanisms, highlighting the “Drop-towards-the-End” phenomenon whereby tokens later in a sequence are at greater risk of being dropped due to experts reaching their maximum capacity limits, an issue that is exacerbated in instruction-tuning datasets. Moreover, OpenMoE identifies a tendency within MoE systems to route tokens based on token-level semantic similarities, leading to “Context-independent Specialization”. Additionally, this token ID routing specialization is established early in pre-training and remains largely fixed, resulting in a consistent pattern of token processing by the same experts throughout training, a phenomenon referred to as “Early Routing Learning”.
@@
Other Advancements on Token-Choice Gating. Despite the implementation of gating heuristics and auxiliary expert-balancing loss functions aimed at achieving a balanced workload distribution among experts, the issue of load imbalance persists as a prevalent challenge within MoE architectures. To solve it, the Balanced Assignment of Sparse Experts (BASE) layer, as conceptualized by Lewis et al. (Lewis et al., 2021) and illustrated in Figure 4 (b), re-envisions the token-to-expert allocation process by casting it as a linear assignment problem, aiming to maximize the token-expert affinities under the constraints that each expert is assigned an equal quantity of tokens. Subsequently, Clark et al. (Clark et al., 2022) introduce a variant of the BASE layer, termed S-BASE, using an optimal transport formulation. Additionally, they devise a reinforcement learning based gating algorithm employing top-1 routing, with the reward function defined as the negative cross-entropy of the predicted token. In addressing the discrete optimization challenge of gating function that can lead to convergence and statistical performance issues when training with gradient-based methods, Hazimeh et al. (Hazimeh et al., 2021) introduce DSelect-k. which is a smooth version of the top- gating algorithm that improves over standard top- gating. This method constitutes a refined version of the top- gating algorithm, featuring enhanced smoothness properties that yield improvements over the conventional top- gating approach. Kudugunta et al. (Kudugunta et al., 2021) diverge from the prevalent token-level gating strategies by introducing a sentence-level gating mechanism. This approach involves generating a sentence representation by averaging the tokens within a sequence and subsequently routing it to an expert. Chi et al.(Chi et al., 2022) observe that prevailing gating mechanisms tend to push hidden representations clustering around expert centroids, implying a trend toward representation collapse, which in turn harms model performance. To counteract this issue, they project hidden vectors into a lower-dimensional space before gating and implement L2 normalization for both token representations and expert embeddings, thus calculating gating scores within a low-dimensional hypersphere. Skywork-MoE (Tianwen Wei, 2024) proposes two innovative techniques: gating logit normalization, which improves expert diversification, and adaptive auxiliary loss coefficients, which provides layer-specific adjustment of auxiliary loss coefficients. Yuan 2.0-M32 (Wu et al., 2024b) proposes a new router network, Attention Router (as illustrated in Figure 4 (e)), which implements a more efficient selection of experts and yields an enhancement in model accuracy over classical linear router network.
@@
Non-trainable Token-Choice Gating. The dynamic training of gating functions within MoE models is standard practice; however, some research has ventured into the realm of non-trainable token-choice gating mechanisms. The most significant benefit of non-trainable token-choice gating is that no additional gating network parameters are required and the full load balancing can be achieved through specific gating mechanisms. The Hash Layer (Roller et al., 2021) utilizes a random fixed gating approach by hashing the input token, achieving competitive results without the necessity of training the gating network. The load balancing is facilitated by the selection of hash functions prior to training, which can equitably distribute token batches. Zuo et al. (Zuo et al., 2021) introduces THOR, an algorithm that randomly allocates two experts to each input during training and inference with a consistency regularized loss promoting consistent predictions. Gururangan et al. (Gururangan et al., 2022) propose the DEMix model, which explicitly assigns distinct experts to discrete pre-training domains, with domain matching being employed to select experts corresponding to the training inputs. Given the potential suboptimality of domain categorization and its limited scope in encompassing test-time domains, a single domain expert selection could undermine the model’s generalizability. To address this, DEMix adopts a parameter-free probabilistic method that dynamically estimates the domain-weighted mixture at inference. Kudugunta et al. (Kudugunta et al., 2021) explore task-level gating incorporating prior knowledge tags, and similarly, M2M-100 model (Fan et al., 2021) utilizes explicit language-specific sublayers with deterministically routing input tokens based on their language. Building upon the aforementioned non-trainable gating strategies—random gating and domain mapping—PanGu- (Ren et al., 2023) presents the Random Routed Experts (RRE) mechanism. As illustrated in Figure 4 (c), this approach initially routes tokens to a domain-specific expert group, followed by a random selection within that group.
@@
In contrast to explicit language-specific expert selection, NLLB (Costa-jussà et al., 2022) leverages trainable gating to manage multilingual machine translation tasks, outperforming the M2M-100 approach (Fan et al., 2021). Addressing task interference in generalist models, Zhu et al. (Zhu et al., 2022) introduce the Conditional MoE, which augments MoE with trainable gating by integrating conditional information at various levels, such as token-level, context-level, modality-level, task-level, and predefined token attributes. Ye et al. (Ye et al., 2022) further investigate the incorporation of trainable gating at task-level MoE. Additionally, STABLEMOE (Dai et al., 2022) identifies a challenge with existing learning-to-route MoE methods: the phenomenon of gating fluctuation. To counter this, STABLEMOE employs a two-stage training process. The first stage focuses on acquiring a balanced and cohesive gating strategy, which is then distilled into a lightweight gate function, decoupled from the backbone model. Subsequently, the second stage leverages the distilled gate for token-to-expert assignments and freezes it to ensure a stable gating strategy throughout further training.
@@
Expert-Choice Gating. Zhou et al. (Zhou et al., 2022a) propose an inversion of the conventional token-choice gating paradigm, wherein each expert selects the top- tokens they will process, as illustrated in Figure 4 (d). This approach circumvents the necessity for auxiliary load balancing losses during training, ensuring a uniform distribution of tokens across experts. However, this method may result in uneven token coverage, with some tokens potentially being processed by multiple experts or not at all. Despite this, the technique demonstrates strong empirical performance and offers an adaptive computational interpretation where the model can implicitly apply more computation to certain tokens. The effectiveness of expert-choice gating is further validated by Zhou et al. in their subsequent Brainformers study (Zhou et al., 2023). Additionally, Komatsuzaki et al. (Komatsuzaki et al., 2022) integrate the expert-choice gating strategy within the Vision Transformer and adapt it for the encoder in T5 models, while maintaining token-choice gating for the T5 decoder.
4.1.2. Dense
@@
In Section 2.1, we discuss the enduring relevance of dense MoE, which activates all the experts for each input process. This dense paradigm continues to inform current innovations in MoE training and inference methodologies, as elaborated in Section 4.4.1. While sparse activation of experts, as a trade-off, may yield efficiency gains at the expense of some performance when compared to a densely activated MoE with an equivalent number of total parameters (Dai et al., 2024; Pan et al., 2024; Shen et al., 2023c), it represents a strategic adjustment to balance computational demands with model capability. Notably, dense activation performs well in the context of LoRA-MoE fine-tuning, where the computational overhead of LoRA experts is comparatively low. This approach enables the effective and flexible integration of multiple LoRAs across a variety of downstream tasks. It preserves the generative capabilities of the original pre-trained model and maintains the unique characteristics of individual LoRAs for each task (Wu et al., 2023b; Dou et al., 2023b).
4.1.3. Soft

### v2

@@ 4.1.1 Sparse
The sparse gating functions activate a selected subset of experts for processing each individual input token, which can be considered as a form of conditional computation [165, 166, 167]. The gating functions have been studied extensively, which may be trained by various forms of reinforcement learning and back-propagation, making binary or sparse and continuous, stochastic or deterministic gating decisions [23, 168, 169, 170, 65]. Shazeer et al. [27] pioneered a differentiable heuristic with auxiliary load balancing losses, in which the outputs from expert computations are weighted by their selection probabilities. This introduces a differentiable aspect to the gating process, thereby facilitating the derivation of gradients that can guide the gating function’s optimization. This paradigm has subsequently become predominant in the realm of MoE research. Due to its selection of experts for each input token, this method can be recognized as a gating function with token choice.
@@
Token-Choice Gating. Shazeer et al. [27] posited the necessity of gating inputs to the top- experts, with , to enhance the efficacy of MoE. The rationale behind this approach is that by simultaneously consulting multiple experts for a given input, the network can effectively weigh and integrate their respective contributions, thereby improving performance. To accommodate the scalability to thousands of experts within a MoE layer, they employ a two-level hierarchical MoE to reduce the branching factor in the context of a large expert count. Subsequent research has largely affirmed that increasing the value of k enhances performance, which has led to the widespread adoption of this top- strategy with . Notwithstanding, the Switch Transformer model [36] has shown that a top-1 gating strategy (as illustrated in Figure 4 (a)) can also yield competitive results, a finding that has been substantiated and adopted by later studies [65]. Furthermore, M6-t [70] proposed a novel variation of the top-1 gating called expert prototyping, which organizes experts into k groups and then applies top-1 gating in each group. Their experimental results show the training and downstream perplexity of a 16-layer model in order of best to worst: expert prototyping with 4 top-1 gating, 1 top-4 gating, 1 top-16 gating, 1 top-1 gating.
@@
Auxiliary Loss for Token-Choice Gating. Token-choice gating algorithms frequently incorporate an auxiliary loss during training to promote equitable token distribution across experts. Table I shows prevalent auxiliary loss functions leveraged in the field. Shazeer et al. [27] quantify the importance of an expert in relation to a training batch via the batchwise sum of the gate values for that expert. They define an additional loss , added to the overall loss function for the model. This loss, which is equal to the square of the coefficient of variation of the set of importance values and multiplied by a hand-tuned scaling factor , encourages all experts to have equal importance. Although promotes balance in importance, it does not guarantee an even distribution of training examples among experts, which can lead to execution inefficiencies on distributed computing environments. To address this, they introduce a second loss to ensure balanced loads. Building on this foundation, GShard [28] define a new differentiable auxiliary loss using a differentiable approximation (the dot-product of mean gates and mean gating decisions per expert), as detailed in Section 2.2. Switch Transformers [36] and many other subsequent studies [35, 29, 31, 68] have embraced this design, and enhancements [69, 33, 67] have been made to cater to diverse requirements. Nevertheless, ST-MoE [37] identified limitations with , particularly at larger scales, leading to unreliable training outcomes. To mitigate this, it introduces the integration of router z-loss , improving training stability without quality degradation by penalizing large logits entering the gating network. Since this loss encourages absolute magnitude of values to be smaller, roundoff errors are reduced, which can be quite impactful for exponential functions such as gating. Additionally, Mod-Squad [71] posits the difficulty of training multi-task models under such an expert-balancing loss, which may inadvertently force experts to set parameters on conflicting tasks or hinder the potential synergies from parameter sharing across complementary tasks. Instead, it proposes to maximize the mutual information (MI) between experts and tasks to build task-expert alignment. Differently, ModuleFormer [73] proposes to maximize the Mutual Information between experts and tokens. Furthermore, DS-MoE [64] extends the application of , calibrating different weightings , in Mixture-of-Attention (MoA, as illustrated in Figure 5 (a)) and FFN MoE modules of different size models.
@@
Expert Capacity for Token-Choice Gating. In conjunction with load balancing via auxiliary loss, GShard [28] incorporates an expert capacity limit, defining a threshold for the number of tokens an expert can process. This can lead to token overflow, where excess tokens are not processed by the designated expert. GShard also proposes a random routing mechanism that selects a secondary expert with a probability proportional to its weight, under the intuition that the contribution of a secondary expert can be negligible, given that the output is a weighted average and the secondary weight is typically small. For the task of image classification with Vision Transformer (ViT) models, Riquelme et al. [8] enhance the top- gating strategy with Batch Prioritized Routing (BPR), which assigns priority based on higher gating scores rather than the sequence order of tokens. Zoph et al. [37] have demonstrated the efficacy of BPR in the context of MoE language models. Kim et al. [76] suggest randomizing token prioritization within sequences to mitigate routing bias towards early-positioned tokens. OpenMoE [38] provides a comprehensive analysis of gating mechanisms, highlighting the “Drop-towards-the-End” phenomenon whereby tokens later in a sequence are at greater risk of being dropped due to experts reaching their maximum capacity limits, an issue that is exacerbated in instruction-tuning datasets. Moreover, OpenMoE identifies a tendency within MoE systems to route tokens based on token-level semantic similarities, leading to “Context-independent Specialization”. Additionally, this token ID routing specialization is established early in pretraining and remains largely fixed, resulting in a consistent pattern of token processing by the same experts throughout training, a phenomenon referred to as “Early Routing Learning”.
@@
Other Advancements on Token-Choice Gating. Despite the implementation of gating heuristics and auxiliary expert-balancing loss functions aimed at achieving a balanced workload distribution among experts, the issue of load imbalance persists as a prevalent challenge within MoE architectures. To solve it, the Balanced Assignment of Sparse Experts (BASE) layer, as conceptualized by Lewis et al. [74] and illustrated in Figure 4 (b), re-envisions the token-to-expert allocation process by casting it as a linear assignment problem, aiming to maximize the token-expert affinities under the constraints that each expert is assigned an equal quantity of tokens. Subsequently, Clark et al. [65] introduce a variant of the BASE layer, termed S-BASE, using an optimal transport formulation. Additionally, they devise a reinforcement learning based gating algorithm employing top-1 routing, with the reward function defined as the negative cross-entropy of the predicted token. The discrete optimization of gating function can lead to convergence and statistical performance issues when training with gradient-based methods. To address these issues, Hazimeh et al. [75] introduce DSelect-k, which is a smooth version of the top- gating algorithm that improves over standard top- gating. This method constitutes a refined version of the top- gating algorithm, featuring enhanced smoothness properties that yield improvements over the conventional top- gating approach. Kudugunta et al. [77] diverge from the prevalent token-level gating strategies by introducing a sentence-level gating mechanism. This approach involves generating a sentence representation by averaging the tokens within a sequence and subsequently routing it to an expert. Chi et al. [80] observe that prevailing gating mechanisms tend to push hidden representations clustering around expert centroids, implying a trend toward representation collapse, which in turn harms model performance. To counteract this issue, they project hidden vectors into a lower-dimensional space before gating and implement L2 normalization for both token representations and expert embeddings, thus calculating gating scores within a low-dimensional hypersphere. Cai et al. [88] identify the difficulty in training the adaptive routers due to gradient vanishing, and introduce a novel strategy involving the training of a Surrogate Model (SM) that predicts LLM’s language loss given only router choices. Once trained, the SM is frozen, and the router is subsequently tuned to minimize language loss, relying exclusively on feedback from the SM. Skywork-MoE [67] proposes two innovative techniques: gating logit normalization, which improves expert diversification, and adaptive auxiliary loss coefficients, which provides layer-specific adjustment of auxiliary loss coefficients. Yuan 2.0-M32 [84] proposes a new router network, Attention Router (as illustrated in Figure 4 (e)), which implements a more efficient selection of experts and yields an enhancement in model accuracy over classical linear router network. Zeng et al. [85] posit that the complexity of token feature abstraction may necessitate a variable number of experts to process. In response, they propose AdaMoE, a novel approach that enables token-adaptive gating for MoE, allowing for a dynamic number of selected experts per token. AdaMoE subtly modifies the standard top- MoE by incorporating a predetermined set of null experts and increasing the value of . Importantly, AdaMoE does not mandate a uniform allocation of null experts across tokens but ensures the average engagement of null experts with a load-balancing loss, resulting in an adaptive number of null/true experts used by each token. Dynamic Mixture of Experts (DYNMoE) [87] also introduces an innovative gating mechanism that enables individual tokens to automatically determine the number of activated experts via the trainable per-expert thresholds, incorporating an adaptive process that automatically add or remove experts during training. Shi et al. [86] introduce Self-Contrast Mixtureof-Experts (SCMoE), a training-free strategy that utilizes the contrastive information among different gating strategies to engage unchosen experts during inference.
@@
Non-trainable Token-Choice Gating. The dynamic training of gating functions within MoE models is standard practice; however, some research has ventured into the realm of non-trainable token-choice gating mechanisms. The most significant benefit of non-trainable token-choice gating is that no additional gating network parameters are required and the full load balancing can be achieved through specific gating mechanisms. The Hash Layer [89] utilizes a random fixed gating approach by hashing the input token, achieving competitive results without the necessity of training the gating network. The load balancing is facilitated by the selection of hash functions prior to training, which can equitably distribute token batches. Zuo et al. [90] introduces THOR, an algorithm that randomly allocates two experts to each input during training and inference with a consistency regularized loss promoting consistent predictions. Gururangan et al. [91] propose the DEMix model, which explicitly assigns distinct experts to discrete pretraining domains, with domain matching being employed to select experts corresponding to the training inputs. Given the potential suboptimality of domain categorization and its limited scope in encompassing test-time domains, a single domain expert selection could undermine the model’s generalizability. To address this, DEMix adopts a parameter-free probabilistic method that dynamically estimates the domain-weighted mixture at inference. Kudugunta et al. [77] explore task-level gating incorporating prior knowledge tags, and similarly, M2M-100 model [92] utilizes explicit language-specific sublayers with deterministically routing input tokens based on their language. Building upon the aforementioned non-trainable gating strategies—random gating and domain mapping—PanGu- [93] presents the Random Routed Experts (RRE) mechanism. As illustrated in Figure 4 (c), this approach initially routes tokens to a domain-specific expert group, followed by a random selection within that group.
@@
In contrast to explicit language-specific expert selection, NLLB [78] leverages trainable gating to manage multilingual machine translation tasks, outperforming the M2M-100 approach [92]. Addressing task interference in generalist models, Zhu et al. [81] introduce the Conditional MoE, which augments MoE with trainable gating by integrating conditional information at various levels, such as token-level, context-level, modality-level, task-level, and predefined token attributes. Ye et al. [79] further investigate the incorporation of trainable gating at task-level MoE. Additionally, STABLEMOE [72] identifies a challenge with existing learning-to-route MoE methods: the phenomenon of gating fluctuation. To counter this, STABLEMOE employs a two-stage training process. The first stage focuses on acquiring a balanced and cohesive gating strategy, which is then distilled into a lightweight gate function, decoupled from the backbone model. Subsequently, the second stage leverages the distilled gate for token-to-expert assignments and freezes it to ensure a stable gating strategy throughout further training.
@@
Expert-Choice Gating. Zhou et al. [94] propose an inversion of the conventional token-choice gating paradigm, wherein each expert selects the top- tokens they will process, as illustrated in Figure 4 (d). This approach circumvents the necessity for auxiliary load balancing losses during training, ensuring a uniform distribution of tokens across experts. However, this method may result in uneven token coverage, with some tokens potentially being processed by multiple experts or not at all. Despite this, the technique demonstrates strong empirical performance and offers an adaptive computational interpretation where the model can implicitly apply more computation to certain tokens. The effectiveness of expert-choice gating is further validated by Zhou et al. in their subsequent Brainformers study [95]. Additionally, Komatsuzaki et al. [49] integrate the expert-choice gating strategy within the Vision Transformer and adapt it for the encoder in T5 models, while maintaining token-choice gating for the T5 decoder.
4.1.2 Dense
@@
In Section 2.1, we discuss the enduring relevance of dense MoE, which activates all the experts for each input process. This dense paradigm continues to inform current innovations in MoE training and inference methodologies, as elaborated in Section 4.4.1. While sparse activation of experts, as a trade-off, may yield computational efficiency gains at the expense of some performance loss when compared to a densely activated MoE with an equivalent number of total parameters [69, 64, 73], it represents a strategic adjustment to balance computational demands with model capability. Notably, dense activation performs well in the context of LoRA-MoE fine-tuning, where the computational overhead of LoRA experts is comparatively low. This approach enables the effective and flexible integration of multiple LoRAs across a variety of downstream tasks. It preserves the generative capabilities of the original pretrained model and maintains the unique characteristics of individual LoRAs for each task [63, 45].
4.1.3 Soft